OpenStudy (anonymous):

What is the equation of the line that passes through the point (2, 3) and is perpendicular to the line y = –x – 3?

OpenStudy (campbell_st):

for the line y = -x -3 the slope(gradient) is \[m _{l} = -1\] a perpendular line has a gradient od \[m _{p} = \frac{-1}{m _{l}}\] \[m _{p} = \frac{-1}{-1} \] gradient of the perpendicular is m = 1 use the point gradient formula m = 1, y1 = 3 and x1 = 2 \[y - y _{1} = m(x -x _{1})\] to find the equation of the line

OpenStudy (anonymous):

these are the choices i have: y = x + 5 y = x + 1 y = –x – 1 y = –x + 5

OpenStudy (campbell_st):

y - 3 = 1(x - 2) y - 3 = x - 2 just simplify add 3 to both sides

OpenStudy (anonymous):

so.. y = x+1 ?